OpenStudy (anonymous):

-7 + x = 13 x = __________

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

20

OpenStudy (uri):

-7 + x = 13 x =13 +7 x =20 ans.

OpenStudy (anonymous):

thanks

OpenStudy (dls):

uri is a born genius

OpenStudy (uri):

o_O